# NeuSemSlice: Towards Effective DNN Model Maintenance via Neuron-level Semantic Slicing

### v1

@@ 2.4. Motivation
The state-of-the-art technique exhibits three noticeable limitations in addressing the challenging tasks of DNN model maintenance: ❶ Lack of granularity. Existing technique is confined to layer level, which lacks the flexibility for optimal model restructuring and adaptation on the finer neuron level. ❷ Ineffectiveness for smaller models. The existing technique is ineffective for smaller DNNs, crucial for on-device deployments, due to their lack of layers with similar semantics. This limits the ability to downsize by removing redundant layers for restructuring or adaptation. ❸ Absence of incremental model development. This crucial aspect of model maintenance is not adequately addressed. The current technique is unable to identify key components at the neuron level (Du et al., 2019), which is essential for the model’s functions, leading to challenges in developing new functions while preserving existing ones with only layer-level adjustments.
@@ 2.5. NeuSemSlice Overview
To address the issues, we propose NeuSemSlice which focuses on neuron-level DNNs semantic decomposition and involves slicing DNNs to identify and isolate critical neuron components, facilitating a variety of model maintenance tasks. Figure 1 shows a detailed workflow of our NeuSemSlice framework, which mainly consists of two modules: Semantic Slicing and Semantic-aware Model Maintenance.
@@
Semantic-aware model maintenance. By utilizing the aggregated critical neurons in the model, this approach supports various applications, offering more robust solutions for long-term DNN model maintenance. Beyond restructuring and re-adapting the model, this method also addresses incremental model development for the first time. It specifically counteracts catastrophic forgetting, which occurs when an AI model learns new information, by stabilizing the critical neurons’ parameters. This enables the model to incrementally develop functions with limited access to the original training data.
@@ 3.1. A Pilot Study on Existing DNNs Slicing Strategies
Global Slicing. One possible strategy is the global approach, which calculates contribution metrics across the entire dataset and then selects the top 10% (i.e., set $k$ in Equation 3 as 10%) of neurons, as indicated by these metrics, as critical neurons. This process, outlined in Section 2.3, results in five distinct sets of critical neurons identified using different metrics. We then devise a strategy to evaluate the importance of each set to determine the most effective contribution metric. Specifically, to assess and compare the significance of the critical neurons identified by various metrics, we adopt the method from Xie et al. (Xie et al., 2022) by zeroing the outputs of these neurons and observing the impact on model predictions. Greater changes in predictions suggest the higher neuron significance.
The study is conducted on CIFAR-10 (Krizhevsky et al., 2009) and training VGG-16 (Simonyan and Zisserman, 2014) for classifying. Table 1 shows that in a global setting, all neuron contribution metrics have limitations. Surprisingly, masking neurons deemed less critical often leads to more significant accuracy losses, suggesting a misalignment between globally calculated contributions and the neurons’ true importance. For example, using the Avg metric, masking 20$\rightarrow$30of neurons has a greater impact than masking the top 10$\rightarrow$20%, as highlighted by the green arrows in the table. The results suggest that global slicing is not effective for model maintenance tasks.
@@
Category-wise Slicing. Another possible strategy is the category-wise approach. Acknowledging that neuron contributions can vary widely between categories, which may limit the effectiveness of global DNN slicing, we conduct a category-specific analysis. We calculate the five metrics for each category and then conduct masking experiments on data from each category (i.e., zeroing the outputs of neurons deemed critical for that category) to assess accuracy on a per-category basis. The averaged and individual category results are shown in Tables 2 and 3, respectively. Masking the top 10% of neurons shows that certain metrics (e.g., Avg and Ens) drastically lowered classification performance, nearly to zero. This demonstrates that a category-specific approach is more effective in identifying critical neurons than that of the global-based apporach.
@@ 3.2. Identified Challenges for Semantic Slicing
Identifying critical neurons for model maintenance involves two main challenges: the optimal proportion of critical neurons varies by category and by layer. ❶ Different category-wise distribution. Integrating category-specific critical neurons into a set for the entire dataset is complex as the ideal proportion of these neurons differs across categories. For example, masking 10% to 20% of critical neurons in Category 4 hardly affects performance, suggesting redundancy. In comparison, masking 30% to 40% in Category 1 significantly impacts performance, indicating their importance for that category. ❷ Different layer-wise distribution. The distribution of critical neurons changes across layers, often becoming sparser in higher layers. Current methods propose a uniform selection range for all categories and layers, which can either overlook crucial neurons or include too many non-critical ones, affecting accuracy and efficiency in model maintenance. Therefore, determining the optimal range of critical neurons for each category and layer is essential for effective DNN slicing.
@@ 3.3. The Design of Semantic Slicing
Problem Definition. Inspired by prior layer similarity evaluation works (Nguyen et al., 2020; Ren et al., 2023), we formally reformulate the task of selecting top $k$ neurons for each layer and category as follows. For each layer $l$ and category $c$, identify $k$ neurons such that they fulfill the condition: $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$, where $\Theta$ represents a unified threshold applicable across various categories and layers, and $\Omega_{k}^{l,c}$ is the selected neuron set. Employing the CKA metric as in Equation 2, we calculate a similarity metric in the range of [0, 1] as follows:
(4) $min|\mathbf{\Omega}^{l,c}|\quad\text{subject to}\quad\frac{M^{\mathbf{\Omega}^{l,c}}\cdot M^{l,c}}{||M^{\mathbf{\Omega}^{l,c}}||\cdot||M^{l,c}||}>\Theta,\vspace{-0.4cm}$
@@
Algorithm Implementation. To address the aforementioned issue, we adopt a Linear Scan strategy as described by Algorithm 1. Specifically, this approach starts with the neurons possessing the top $i$ highest contribution scores in the array $\mathbf{C}^{l,c}$. We initially set $\mathbf{\Omega}^{l,c}$ to encompass these top $i$ neurons and proceed to calculate the similarity score. For each neuron contribution score within $\mathbf{C}^{l,c}$, the strategy involves sequentially selecting the top $i$ neurons with the highest contribution scores from the neurons not in $\mathbf{\Omega}^{l}$ (i.e., $\mathbf{N}^{l}\setminus\mathbf{\Omega}^{l,c}$). This process is continued until the condition $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$ is met.
To reduce the computational burden involved in incrementally adding neurons to the subset and computing their semantic similarity with all neurons in the layer, we can increase the value of $i$.
@@
Result: The critical neuron set for category $c$: $\mathbf{\Omega}^{l,c}$.
1 $\Omega^{l,c}\leftarrow\emptyset$
2while $\mathrm{sim}_{\mathrm{sem}}(f^{\Omega^{l,c}},f^{l,c})<\Theta$ do
@@ 1 $\Omega^{l,c}\leftarrow\emptyset$
Categorical Critical Neurons Aggregation. After determining the critical neurons for each category, they can be combined on a category basis to establish the critical neuron set for the entire model, as shown in Figure 1. For an entire dataset $D$ containing $m$ categories, this can be expressed as: $\mathbf{\Omega}^{l,D}=\mathbf{\Omega}^{l,D_{c_{0}}}\cup\mathbf{\Omega}^{l,D_{c_{1}}}\cup\ldots\cup\mathbf{\Omega}^{l,D_{c_{m-1}}}$. We can further depict the neuron contribution of the model $f$ as $\mathbf{\Omega}=\{\mathbf{\Omega}^{l_{0,D}},\mathbf{\Omega}^{l_{1,D}},...,\mathbf{\Omega}^{l_{L-1,D}}\}$. Given the neuron contribution metrics $\mathbf{C}$, the selection threshold $\Theta$ directly determines $\mathbf{\Omega}$, simplifying the process by eliminating the need to consider varying optimal neuron proportions across categories and layers.
@@ 4.1. Model Restructure
$I(n)=\left\{\begin{array}[]{ll}1&\text{if }n\in\mathbf{\Omega},\\ 0&\text{otherwise}.\end{array}\right.$
Then each neuron $n^{\prime}$ in the restructured model $f^{\prime}$ could be expressed as $n^{\prime}=n\cdot I(n)$. Note that the threshold value $\Theta$ can be dynamically adjusted to balance model performance and size during restructuring. By adjusting $\Theta$, this method allows for a restructured model that varies in accuracy and size to suit diverse needs. Such restructuring will lead to an effective decrease in the storage space. Moreover, some weight-level unstructured pruning techniques, such as (Lee et al., 2020; Morcos et al., 2019; Gale et al., 2019; Evci et al., 2019), can be combined with our method. By leveraging model restructure, these pruning methods can focus on redundant weights within critical neurons, thus accelerating the pruning process and further reducing the model size.
@@
4.2. Model Re-adaption
Model re-adaptation improves adaptability and robustness by focusing on training parameters related to critical neurons. Previous research (Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021) has shown that in cases of poor performance, such as adversarial examples, adversarial noise is amplified and propagated through critical neurons. Therefore, the fundamental cause is the abnormal behavior of critical neurons leading to errors, while other neurons are less affected. Retraining these critical neurons’ parameters can thus more efficiently repair and enhance the model in terms of cases with poor performance. The complete model re-adaption workflow is shown in Figure 3. Using adversarial samples as an example, the retraining strategy adopts the following loss function:
$\mathcal{L}=\mathcal{L}_{c}(x,y;\theta)+\mathcal{L}_{c}(adv_{x},y;\theta),$
where $L_{c}$ denotes the cross entropy, $adv_{x}$ is the adversarial counterpart of the sample $x$, and the notation $\theta$ represents the parameters in model $f$. The parameters update strategy in the retraining is:
$\theta^{\prime}_{\mathbf{\Omega}}=\theta_{\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{\Omega}}}\mathcal{L},$
where $\eta$ is the learning rate. Note that in this strategy, the parameters that fall outside the critical neurons set $\mathbf{\Omega}$ remain unchanged and are not updated.
Figure 3. Model Re-adaption based on semantic slicing. We train only the critical neurons for model re-adaptation.
@@ 4.3. Model Incremental Development
The model incremental development refers to the DNN model to learn from new data over time without forgetting previously learned information. This is crucial for creating DNN models that can adapt and evolve in dynamic environments. Taking an image classification designed to recognize different types of animals as an example, we initially only have data to train a DNN to classify images into three categories: Dogs, Cats, and Birds (old task). After the initial training, we collect more data and want the DNN to learn to recognize two new categories Fish and Horses (new task). It is a significant challenge for DNN models. ❶ In practical applications, the training data for the old task might be discarded to save storage space and we only have access to the data of the new task. When training only with the data of the new task, the incremental development of DNNs is prone to ‘catastrophic forgetting’, where learning new tasks can lead to the loss of old tasks. ❷ Even with the training data of the old task, the training process is less resource-efficient because the DNN models have to be retrained with the entire dataset (new and old data).
@@
This issue is particularly prominent in real-world scenarios, such as in IoT devices (Liu et al., 2021), where incremental development is essential. Due to space constraints, these devices might initially be trained on a limited set of data categories. However, as new data becomes available, it is often impractical to retrain the entire model from scratch or to delete previously learned categories.
Our strategy achieves model incremental development by freezing critical neurons’ weights during new task learning, preserving performance on original tasks while allowing new skills acquisition by training the model’s ‘redundant’ components. Specifically, for a model that has already been trained for task $A$, our initial step involves identifying the critical neurons associated with task $A$. Subsequently, when training for a new task, task $B$, we fix the parameters related to these critical neurons as shown in Figure 4. This approach allows the model to learn $B$ by focusing on training the non-critical neurons. Since the critical neurons and their parameter settings preserve the critical information of task $A$, we maintain the predictive accuracy for task $A$ by retaining the critical neurons.
The model incremental development strategy follows the following loss:
$\mathcal{L}=\mathcal{L}_{c}(x_{B},y_{B};\theta),$
where $L_{c}$ denotes the cross entropy and $(x_{B},y_{B})$ represents data pairs from task B. The parameters update strategy is:
@@
$\theta^{\prime}_{\mathbf{N}\setminus\mathbf{\Omega}}=\theta_{\mathbf{N}\setminus\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{N}\setminus\mathbf{\Omega}}}\mathcal{L},$
@@
where $\eta$ is the learning rate, $\mathbf{N}$ represents all neurons in model $f$, and $\mathbf{N}\setminus\mathbf{\Omega}$ represents the neurons outside the critical neurons set $\mathbf{\Omega}$. In the inference process, we will exclusively utilize the frozen parameters associated with task $A$ for inference on task A, and employ the remaining parameters for task B.

### v2

@@ 2.4. Motivation
Maintaining DNN models presents significant challenges, especially as both models and datasets continue to scale in size and complexity. Current state-of-the-art techniques (Ren et al., 2023) exhibit several limitations that impede effective model maintenance:
Insufficient Granularity. Existing methods predominantly operate at the layer level, which restricts their ability to perform precise model restructuring and adaptation. This coarse granularity overlooks critical details at the neuron level, where fine-grained adjustments to individual neurons—especially those supporting specific functionalities—can significantly enhance model performance and efficiency.
Ineffectiveness for Relatively Small Models. The term “small models” in the context of this paper often refers to models that are relatively small compared to the size of their training datasets, rather than being absolutely small. DeepArc (Ren et al., 2023) has demonstrated that layer-level modularization is most effective when the model’s scale significantly exceeds that of the dataset. For instance, DeepArc’s experiments utilized large models such as ResNet-110, Wide-ResNet-38, and VGGNet-19 trained on relatively small datasets like MNIST, FMNIST, and CIFAR-10. Our own experiments (RQ1) further confirm that when the disparity between model size and dataset size is reduced, the benefits of layer-level modularization diminish or disappear entirely. This indicates that layer-level maintenance techniques are less effective for models that are not excessively large relative to their training data. Consequently, there is a need for more refined, neuron-level approaches to ensure efficient and scalable model maintenance across varying model and dataset scales.
Lack of Incremental Model Development. Incremental model development is a crucial aspect of model maintenance that remains inadequately addressed by current techniques. Existing approaches struggle to identify and preserve key components at the neuron level (Du et al., 2019), which are essential for maintaining the model’s functionality. This limitation poses significant challenges in developing new features or functions while ensuring that existing capabilities are retained, especially when relying solely on layer-level adjustments. A neuron-level maintenance strategy would enable more precise and incremental updates, facilitating the continuous evolution of DNN models without compromising their existing performance.
Addressing these limitations requires a shift towards more granular maintenance methodologies. By adopting neuron-level strategies, it becomes possible to achieve finer control over model restructuring and adaptation, enhance the effectiveness of maintenance for models of varying sizes relative to their datasets, and support incremental development processes. This approach promises to provide a more flexible and robust framework for DNN model maintenance, ultimately leading to more efficient and scalable deployment of deep learning models in diverse applications within the field of software engineering.
@@ 2.5. NeuSemSlice Overview
To address the issues, we propose NeuSemSlice which focuses on neuron-level DNN semantic decomposition and involves slicing DNNs to identify and isolate critical neuron components, facilitating a variety of model maintenance tasks. Figure 1 shows a detailed workflow of our NeuSemSlice framework, which mainly consists of two modules: Semantic Slicing and Semantic-aware Model Maintenance.
@@
Semantic-aware model maintenance. By utilizing the aggregated critical neurons in the model, this approach supports various applications, offering more robust solutions for long-term DNN model maintenance. Beyond restructuring and re-adapting the model, this method also addresses incremental model development. It specifically counteracts catastrophic forgetting, which occurs when an AI model learns new information, by stabilizing the critical neurons’ parameters. This enables the model to incrementally develop functions with limited access to the original training data.
@@ 3.1. A Pilot Study on Existing DNNs Slicing Strategies
Global Slicing. One possible strategy is the global approach, which calculates contribution metrics across the entire dataset and ranks all neurons accordingly. This process, outlined in Section 2.3, results in five distinct sets of critical neurons identified using different metrics. We then group neurons into different percentile ranges (e.g., top 0–10%, 10–20%, etc.) and devise a strategy to evaluate the importance of each set to determine the most effective contribution metric. Specifically, to assess and compare the significance of the critical neurons identified by various metrics, we adopt the method from Xie et al. (Xie et al., 2022) by zeroing the outputs of these neurons and observing the impact on model predictions. Greater changes in predictions suggest higher neuron significance.
The study is conducted on CIFAR-10 (Krizhevsky et al., 2009), using VGG-16 (Simonyan and Zisserman, 2015) for classifying. Table 1 shows that in a global setting, all neuron contribution metrics have limitations. Surprisingly, masking neurons deemed less critical often leads to more significant accuracy losses, suggesting a misalignment between globally calculated contributions and the neurons’ true importance. For example, using the Avg metric, masking 20$\rightarrow$30of neurons has a greater impact than masking the top 10$\rightarrow$20%, as highlighted by the green arrows in the table. The results suggest that global slicing is not effective for model maintenance tasks.
@@
Category-wise Slicing. Another possible strategy is the category-wise approach. Acknowledging that neuron contributions can vary widely between categories, which may limit the effectiveness of global DNN slicing, we conduct a category-specific analysis. We calculate the five metrics for each category and then conduct masking experiments on data from each category (i.e., zeroing the outputs of neurons deemed critical for that category) to assess accuracy on a per-category basis. The averaged and individual category results are shown in Tables 2 and 3, respectively. Masking the top 10% of neurons shows that certain metrics (e.g., Avg and Ens) drastically lowered classification performance, nearly to zero. This demonstrates that a category-specific approach is more effective in identifying critical neurons than that of the global-based approach.
@@ 3.2. Identified Challenges for Semantic Slicing
Identifying critical neurons for model maintenance involves two main challenges: the optimal proportion of critical neurons varies by category and by layer. ❶ Different category-wise distribution. Integrating category-specific critical neurons into a set for the entire dataset is complex as the ideal proportion of these neurons differs across categories. For example, masking 10% to 20% of critical neurons in Category 4 hardly affects performance, suggesting redundancy. In comparison, masking 30% to 40% in Category 1 significantly impacts performance, indicating their importance for that category, as shown in Table 3. ❷ Different layer-wise distribution. The distribution of critical neurons changes across layers, often becoming sparser in higher layers. Current methods propose a uniform selection range for all categories and layers, which can either overlook crucial neurons or include too many non-critical ones, affecting accuracy and efficiency in model maintenance. Therefore, determining the optimal range of critical neurons for each category and layer is essential for effective DNN slicing.
@@ 3.3. The Design of Semantic Slicing
Problem Definition. Inspired by prior layer similarity evaluation works (Nguyen et al., 2021; Ren et al., 2023), we formally reformulate the task of selecting top $k$ neurons for each layer and category as follows. For each layer $l$ and category $c$, identify $k$ neurons such that they fulfill the condition: $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$, where $\Theta$ represents a unified threshold applicable across various categories and layers, and $\Omega_{k}^{l,c}$ is the selected neuron set. Employing the CKA metric as in Equation 2, we calculate a similarity metric in the range of [0, 1] as follows:
(4) $min|\mathbf{\Omega}^{l,c}|\quad\text{subject to}\quad\frac{M^{\mathbf{\Omega}^{l,c}}\cdot M^{l,c}}{||M^{\mathbf{\Omega}^{l,c}}||\cdot||M^{l,c}||}>\Theta,\vskip-11.38092pt$
@@
Algorithm Implementation. To address the aforementioned issue, we adopt a Linear Scan strategy as described by Algorithm 1. We break down the implementation into three steps:
  1. (1)
    Initialization: We begin by selecting the top $i$ neurons with the highest contribution scores from $\mathbf{C}^{l,c}$ to form the initial set $\mathbf{\Omega}^{l,c}$. At this stage, $\mathbf{\Omega}^{l,c}$ contains a small but promising subset of neurons for category $c$ at layer $l$. Using this initial set, we compute $\mathrm{sim}_{\mathrm{sem}}\bigl{(}f^{\mathbf{\Omega}^{l,c}},f^{l,c}\bigr{)}$ to assess how well it captures the semantic characteristics of the entire neuron population $\mathbf{N}^{l}$.
  2. (2)
    Iteration Until Threshold: If the current similarity does not surpass the predefined threshold $\Theta$, we continue the selection process. This involves identifying the next top $i$ neurons from the remaining pool $\mathbf{N}^{l}\setminus\mathbf{\Omega}^{l,c}$ according to their contribution scores. By iteratively adding these high-contribution neurons, we gradually refine the subset $\mathbf{\Omega}^{l,c}$, ensuring that it becomes more semantically aligned with $f^{l,c}$.
  3. (3)
    Stopping Condition: The iteration proceeds until we achieve $\mathrm{sim}_{\mathrm{sem}}\bigl{(}f^{\mathbf{\Omega}^{l,c}},f^{l,c}\bigr{)}>\Theta$. Once the threshold is met, the algorithm terminates, and the resulting set $\mathbf{\Omega}^{l,c}$ is deemed critical for capturing the semantic features of category $c$ at layer $l$. To further reduce the computational load, practitioners may choose a larger increment $i$, thereby requiring fewer iterations and faster convergence.
@@
Result: The critical neuron set for category $c$ at layer $l$: $\mathbf{\Omega}^{l,c}$.
1
[…]
3while $\mathrm{sim}_{\mathrm{sem}}\bigl{(}f^{\Omega^{l,c}},f^{l,c}\bigr{)}<\Theta$ do
@@
Categorical Critical Neurons Aggregation. After determining the critical neurons for each category, they can be combined on a category basis to establish the critical neuron set for the entire model, as shown in Figure 1. For an entire dataset $D$ containing $m$ categories, this can be expressed as: $\mathbf{\Omega}^{l,D}=\mathbf{\Omega}^{l,D_{c_{0}}}\cup\mathbf{\Omega}^{l,D_{c_{1}}}\cup\ldots\cup\mathbf{\Omega}^{l,D_{c_{m-1}}}$. We can further represent the critical neuron set of the model $f$ as $\mathbf{\Omega}=\{\mathbf{\Omega}^{l_{0,D}},\mathbf{\Omega}^{l_{1,D}},...,\mathbf{\Omega}^{l_{L-1,D}}\}$. Given the neuron contribution metrics $\mathbf{C}$, the selection threshold $\Theta$ directly determines $\mathbf{\Omega}$, simplifying the process by eliminating the need to consider varying optimal neuron proportions across categories and layers.
@@ 4.1. Model Restructure
(7) $I(n)=\left\{\begin{array}[]{ll}1&\text{if }n\in\mathbf{\Omega},\\ 0&\text{otherwise}.\end{array}\right.$
Then each neuron $n^{\prime}$ in the restructured model $f^{\prime}$ could be expressed as $n^{\prime}=n\cdot I(n)$. Note that the threshold value $\Theta$ can be dynamically adjusted to balance model performance and size during restructuring. By adjusting $\Theta$, this method allows for a restructured model that varies in accuracy and size to suit diverse needs. Such restructuring will lead to an effective decrease in the storage space.
Moreover, some existing pruning techniques (e.g., (Lee et al., 2021; Morcos et al., 2019; Gale et al., 2019; Evci et al., 2020)) primarily operate at the weight level using unstructured approaches. These methods rely on approximate calculations to reduce weights while maintaining approximate model performance. In contrast, our proposed model restructure method identifies semantic slices (i.e., critical neurons) at the neuron level and removes non-critical neurons to compress the model. By leveraging model restructure, our method effectively complements existing pruning techniques, enabling these weight-based pruning methods to focus on redundant weights within critical neurons, thereby accelerating the pruning process and achieving better model compression.
@@
4.2. Model Re-adaptation
Model re-adaptation improves adaptability and robustness by focusing on training parameters related to critical neurons. Previous research (Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021) has shown that in cases of poor performance, such as adversarial examples, adversarial noise is amplified and propagated through critical neurons. Therefore, the fundamental cause is the abnormal behavior of critical neurons leading to errors, while other neurons are less affected. Retraining these critical neurons’ parameters can thus more efficiently repair and enhance the model in terms of cases with poor performance. The complete model re-adaptation workflow is shown in Figure 3.
[…]
(8) $\mathcal{L}=\mathcal{L}_{c}(x,y;\theta)+\mathcal{L}_{c}(adv_{x},y;\theta),$
where $L_{c}$ denotes the cross entropy, $adv_{x}$ is the adversarial counterpart of the sample $x$, and the notation $\theta$ represents the parameters in model $f$. By summing the losses from both clean and adversarial samples, this objective ensures that the critical neurons learn to correctly classify benign inputs while simultaneously increasing their resistance to adversarial noise. The inclusion of adversarial data in the objective effectively provides stronger gradient signals focused on those regions of the parameter space associated with vulnerability, guiding critical neurons to adjust their representations.
To implement this targeted retraining, the parameter update strategy is carefully restricted. Instead of applying gradient-based updates to all parameters in the model, the procedure focuses exclusively on the parameters associated with the critical neuron set $\mathbf{\Omega}$. Formally, given a learning rate $\eta$, the parameter update rule is:
(9) $\theta^{\prime}_{\mathbf{\Omega}}=\theta_{\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{\Omega}}}\mathcal{L},$
This selective update confines optimization to parameters identified as critical, effectively targeting the root causes of performance issues.
Figure 3. Model Re-Adaptation Based on Semantic Slicing. We Train Only the Critical Neurons for Model Re-Adaptation.
@@ 4.3. Model Incremental Development
Model incremental development refers to the ability of a DNN model to learn from new data over time without forgetting previously learned information. This is crucial for creating DNN models that can adapt and evolve in dynamic environments. Taking an image classification designed to recognize different types of animals as an example, we initially only have data to train a DNN to classify images into three categories: Dogs, Cats, and Birds (old task). After the initial training, we collect more data and want the DNN to learn to recognize two new categories Fish and Horses (new task). This poses a significant challenge for DNN models: ❶ In practical applications, the training data for the old task might be discarded to save storage space and we only have access to the data of the new task. When training only with the data of the new task, the incremental development of DNNs is prone to “catastrophic forgetting”, where learning new tasks can lead to the loss of old tasks. ❷ Even with the training data of the old task, the training process is less resource-efficient because the DNN models have to be retrained with the entire dataset (new and old data).
@@
This issue is particularly prominent in real-world scenarios, such as in Internet of Things (IoT) devices (Liu et al., 2021), where incremental development is essential. Due to space constraints, these devices might initially be trained on a limited set of data categories. However, as new data becomes available, it is often impractical to retrain the entire model from scratch or to delete previously learned categories.
Our strategy enables model incremental development by leveraging the identified critical neurons as stable anchors of previously acquired knowledge. By freezing the weights of these critical neurons during subsequent training phases, the model retains the essential semantic features learned for earlier tasks, thus mitigating catastrophic forgetting. At the same time, the non-critical neurons remain free to adapt, allowing the model to integrate new skills associated with a novel task without diluting its mastery of older tasks.
Concretely, suppose the model has been previously trained on task $A$. We begin by determining the critical neuron set $\mathbf{\Omega}$ that captures task $A$’s essential semantics. These critical neurons, extracted through our earlier semantic slicing procedure, form a robust internal representation of the concepts necessary for solving task $A$. When the model encounters a new task $B$, it relies on these preserved, critical neurons to maintain performance on task $A$ while directing the training process for task $B$ to the non-critical neurons (i.e., those not in $\mathbf{\Omega}$). Figure 4 illustrates this idea: the critical neurons, having proven essential for the original task, remain fixed, while the rest of the network updates to learn the new task.
This approach ensures that the parameters governing the original task’s key semantic representations remain intact, thus safeguarding previously learned knowledge. The model’s representational capacity outside $\mathbf{\Omega}$ is repurposed for incremental learning, enabling it to acquire new skills with minimal interference to existing knowledge. Through this selective updating mechanism, the model incrementally expands its repertoire of competencies in a stable and controlled manner.
Formally, for training on the new task $B$, we adopt the following loss function:
(10) $\mathcal{L}=\mathcal{L}_{c}(x_{B},y_{B};\theta),$
where $\mathcal{L}_{c}$ denotes the cross-entropy loss and $(x_{B},y_{B})$ are the training samples and labels from task $B$. Given this objective, the parameter update rule becomes:
@@
(11) $\theta^{\prime}_{\mathbf{N}\setminus\mathbf{\Omega}}=\theta_{\mathbf{N}\setminus\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{N}\setminus\mathbf{\Omega}}}\mathcal{L},$
@@
where $\eta$ is the learning rate, $\mathbf{N}$ represents the full set of neurons, and $\mathbf{N}\setminus\mathbf{\Omega}$ is the set of neurons not identified as critical for task $A$.
In this manner, only the parameters not involved in encoding the critical semantics from task $A$ are adjusted, allowing the model to assimilate the new task $B$ without eroding the prior knowledge. During inference, the model employs the frozen, critical parameters for predictions on task $A$, ensuring stable performance on previously learned content. For task $B$, it leverages the updated subset of parameters dedicated to the new skill, allowing the model to switch between tasks while maintaining overall system integrity and performance.